The vapor pressure of an aqueous
solution is found to be 24.90 mmHg
at 25 oC. What is the mole fraction
of solute in this solution? The vapor
pressure of water is 25.756 mmHg
at 25oC.
1
Expert's answer
2020-04-22T09:19:01-0400

Raoult's law: (p0 - p)/p0 = x

here p0 is water vapor pressure, p is solution vapor pressure, x is mole fraction of solute.

Mole fraction of solute is:

x = (25.756 - 24.90)/25.756 = 0.033 or 3.3 %